### v2

# Dynamic Strategy Adaptation in Multi-Agent Environments with Large Language Models

Abstract
Large language models (LLMs) demonstrate strong reasoning abilities across mathematical, strategic, and linguistic tasks, yet little is known about how well they reason in dynamic, real-time, multi-agent scenarios, such as collaborative environments in which agents continuously adapt to each other’s behavior, as in cooperative gameplay settings. In this paper, we bridge this gap by combining LLM-driven agents with strategic reasoning and real-time adaptation in cooperative, multi-agent environments grounded in game-theoretic principles such as belief consistency and Nash equilibrium. The proposed framework applies broadly to dynamic scenarios in which agents coordinate, communicate, and make decisions in response to continuously changing conditions. We provide real-time strategy refinement and adaptive feedback mechanisms that enable agents to dynamically adjust policies based on immediate contextual interactions, in contrast to previous efforts that evaluate LLM capabilities in static or turn-based settings. Empirical results show that our method achieves up to a 26% improvement in return over PPO baselines in high-noise environments, while maintaining real-time latency under 1.05 milliseconds. Our approach improves collaboration efficiency, task completion rates, and flexibility, illustrating that game-theoretic guidance integrated with real-time feedback enhances LLM performance, ultimately fostering more resilient and flexible strategic multi-agent systems.
1 Introduction
Large language models (LLMs) are an effective machine learning tool that can perform a broad range of reasoning tasks, such as mathematical problem solving, code generation, commonsense inference, and strategic board games. However, their ability to operate effectively in real-time, multi-agent coordination settings remains under-explored [1]. Given the rise of multi-agent systems, characterized by dynamic collaboration and automation, there is a need for new frameworks capable of supporting real-time strategic adaptation. Although game theory offers a robust mathematical framework for modeling strategic interactions, it typically assumes fixed decision rules that lack the flexibility to accommodate real-world complexity and change [1]. In contrast, reinforcement learning (RL) provides a flexible mechanism for learning through experience; however, it requires extensive training and often produces strategies that lack clear interpretability. To guide this investigation, we ask the following research question: How can LLMs dynamically adapt and refine their game strategies based on real-time feedback in multi-agent, multi-turn environments to achieve enhanced rationality and equilibrium outcomes?
A number of prior works have [2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13] studied negotiation, tool integration, real-time planning, and embodied reasoning in the context of LLM-based decision-making. For example, He et al. [2] and Lewis et al. [3] explore end-to-end learning for negotiation dialogues, while Toolformer [4] demonstrates how LLMs can autonomously learn to invoke external tools during generation. PaLM-E [5] highlights the value of integrating visual and proprioceptive inputs for grounded multimodal reasoning. ReAct prompting [6] illustrates how models can reason and act by interleaving tool calls with intermediate steps. NegotiationArena [7] and StarCraft II benchmarks [8] provide multi-agent coordination evaluations, and CICERO [9] showcases strategic interaction through dialogue in high-stakes, turn-based environments. Recent frameworks such as CAMEL [10], AutoGen [11], LangAgent [12], and Voyager [13] have introduced collaborative LLM teams with structured roles and planning objectives.
Other solutions such as SWE-agent [14] and MetaGPT [15] propose programming-based role assignment and skill composition pipelines, yet often operate with fixed dialogue states and precomputed coordination graphs. OctoAgents [16] explore zero-shot LLM team formation but require architectural homogeneity and lack continual learning mechanisms. The MindAct [17] and CoLLM [18] models examine collaboration over temporal chains but still assume a known world model. While prior solutions [2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13] enable symbolic reasoning, tool usage, and basic interagent dialogue, they typically do not support continuous strategic adaptation in cooperative, real-time environments. Most of these existing approaches operate in static settings or rely on single-shot policy inference. There is little ability for agents to revise goals, respond to degraded teammate performance, or restructure plans in response to unforeseen environment transitions. A central open challenge is, thus, how to enable agents to adjust their strategy mid-execution using context-aware feedback, memory persistence, symbolic priors, and feedback-aware control loops that reflect real-time game-theoretic dynamics.
An examination of the literature that addresses uncertainty, partial observability, and strategic adaptation in games reveals several key limitations [19, 20, 21, 13, 22, 23, 24, 9, 25]. GTBench [19] tests LLMs on game-theoretic matrix tasks, finding poor robustness in partially observable environments. GameTheoryLLM [20] formalizes Bayesian agent workflows for negotiation but imposes synchronous update protocols. RationalLLM [21] studies equilibrium behavior but does not support belief evolution or policy trajectory tracking. MARL benchmarks like MAgent [26] and EnvAgent [27] test emergent roles but fail to align agents with structured symbolic constraints. Systems like MindAgent [28], DEAM [29], and ToM-GPT [30] simulate theory-of-mind and belief graphs but rely on heavily curated scenarios. CoaLLM [31] and OpenAgents [32] attempt to unify self-play and LLM planning, though they still assume pre-scripted task flows.
Although CICERO [9] offers a rich symbolic policy in the context of Diplomacy, it is limited to turn-based play. Reinforcement learning from human feedback (RLHF) [25] allows reward shaping but lacks dynamic revision based on environment evolution. Existing frameworks such as like RolePlay [33] and ThoughtAgents [34] aim to inject personality and dialogue abstraction but do not connect memory or belief state with performance feedback. Few systems track and utilize contextual priors across time steps. Even fewer are capable of aligning their behavioral revisions with task-level metrics like regret, coordination efficiency, or Nash deviation. We argue that bridging this gap requires dynamic feedback systems that are tightly coupled with symbolic modeling and game-theoretic planning principles. The field still lacks a unified method for enabling LLM-based agents to revise their strategies in real-time as goals shift, agents deviate, or coordination structures deteriorate.
The strategic use of LLMs in dynamic, real-time, multi-agent environments faces a significant open problem: how to enable agents to reason adaptively, coordinate under uncertainty, and revise decisions as environments evolve. While early systems [19, 20, 21, 13, 22, 23, 9] have made progress in specific areas, such as static equilibrium modeling, symbolic dialogue protocols, or role-based planning, they remain limited in three key aspects: 1) they are unable to revise beliefs in response to shifting goals or noisy conditions, 2) they lack real-time mechanisms for coordination across agents, and 3) they do not incorporate symbolic priors or game-theoretic principles into feedback loops. These limitations highlight the need for integrated multi-agent reasoning frameworks that support real-time strategy adaptation and structured coordination, guided by symbolic feedback and contextual understanding under uncertainty.
The main contribution of this paper is to address the limitations of static decision-making methods, including rule-based strategies, turn-based reasoning, and offline reinforcement learning, by proposing a new framework that enables agents to adapt strategies in real time using LLM-guided reward shaping. By integrating symbolic evaluations from a frozen language model with a reinforcement learning loop, the proposed system refines agent behavior based on contextual signals such as order progress and timestep. This hybrid framework is designed to support collaboration in unpredictable, noisy environments where conventional policies tend to fail. In addition to developing the architecture, the proposed framework is empirically evaluated across dynamic environments to assess performance, robustness to noise, and latency tradeoffs. The proposed method delivers strategic benefits while preserving real-time inference speeds. In summary, the key contributions of this paper are:
  • We assess the performance of our framework over the Overcooked-AI [35] simulation environment. In these simulations, across five dynamic, partially observable, and high-noise environments, we design and implement a real-time shaping framework that connects proximal policy optimization (PPO)-based agents with frozen LLM evaluations. By leveraging prompt-based symbolic feedback during execution, the framework enables agents to revise decisions mid-episode in response to uncertainty, teammate variability, and timing disruptions. Agents generate structured prompts that encode state, goals, and local history, and use returned judgments to shape the reward signal in a latency-constrained control loop.
  • We formalize symbolic feedback in multi-agent coordination by adapting game-theoretic principles such as belief revision, strategic consistency, and role alignment into a reinforcement learning-compatible reward-shaping pipeline. This symbolic layer guides agents toward cooperative behaviors by enforcing priors on mutual belief, future intent, and role commitment, without requiring full observability or centralized policy access.
  • Across five randomized, high-noise environments, including joint observation noise and delayed timing inputs, the results show that the proposed PPO-LLM agents achieved up to a 26% improvement in average episodic return compared to PPO-only baselines. These gains were robust across different seeds and included rare successful completions under adversarial disruptions, where baseline agents consistently failed.
  • In terms of real-time performance, PPO-LLM agents operated with an average per-action latency of 1.03 to 1.05 milliseconds, inclusive of prompt generation and LLM inference. This satisfies Overcooked-AI’s latency thresholds and confirms the framework’s suitability for deployment in real-time multi-agent infrastructures with frequent decision cycles.
  • PPO-LLM agents also demonstrated a 35% higher full-task completion rate under noise, reflecting more reliable long-horizon coordination and better adaptation to mid-task disruptions. This includes recovery from failed pickups, bottlenecked transitions, and role misassignments under noisy state transitions.
  • The proposed approach allowed the agents to achieve a 22.7% reduction in average idle steps across evaluation episodes, indicating smoother temporal coordination, less waiting, and tighter synchronization between role-dependent sub-tasks. This reflects better implicit communication and increased responsiveness to evolving task conditions.
  • Finally, PPO-LLM agents consistently converged to Nash-like equilibria in cooperative settings such as the Burger Kitchen scenario. Our results in this scenario show that the agents will adopt stable joint policies where deviations became suboptimal. This emergent equilibrium was achieved using symbolic feedback alone, without explicit joint training, parameter sharing, or centralized supervision.
The rest of this paper is organized as follows. Section 2 presents the system architecture, including the real-time shaping pipeline, symbolic feedback loop, and PPO policy integration. Section 3 provides empirical results from Overcooked-AI environments, including performance under noise, task completion rate, latency tradeoffs, and convergence to equilibrium. Section 4 concludes with a summary of findings, limitations, and directions for future research on real-time strategic adaptation using language-based feedback.
2 System Architecture
Figure 1: Real-Time Shaping System with PPO + Frozen LLM. The environment emits a grid-state at each timestep. That grid is flattened into a 1D observation vector and passed to both the PPO agent and a prompt generator. In parallel, a prompt generator encodes task context (e.g., remaining orders and step count) into a text prompt. This prompt is evaluated by a frozen LLM (GPT-Neo 1.3B), whose binary "good"/"bad" signal is mapped to a shaping bonus via the reward-shaping module. The PPO agent then updates its policy using this shaped reward signal.
To support dynamic strategy adaptation in cooperative multi-agent environments, we introduce a real-time reward shaping architecture that integrates PPO with feedback from a frozen LLM. This system is designed to enhance agent coordination through symbolic reasoning and structured evaluations, without modifying the LLM weights or retraining during execution. It consists of six interconnected modules that operate simultaneously during each timestep: i) environment state emitter, ii) observation preprocessor, iii) PPO policy network, iv) prompt generation module, v) frozen LLM evaluator, and vi) reward shaping unit, as shown in Figure 1.
At time $t$, the environment emits a grid state $s_{t}$, which represents the spatial layout of $N=2$ agents, interactive objects (such as ingredients and plates), and task-relevant regions (such as delivery counters). Each grid cell encodes the type and state of any entity occupying it. This state is transformed into a one-dimensional observation vector to ensure compatibility with both the PPO policy network and the LLM prompt generator. We instantiate this structure using the Overcooked-AI framework [35], a cooperative multi-agent simulator in which agents must coordinate to cook and deliver meals under time pressure and spatial constraints. Figure 2 illustrates our specific “Burger Kitchen” configuration, where one agent prepares cooked patties while the other assembles and delivers finished burgers. This layout introduces strict temporal dependencies and spatial bottlenecks that require dynamic role allocation and real-time strategic coordination.
The PPO agent, implemented using Stable-Baselines3 [36], takes the observation and computes an action $a_{t}$ based on its learned policy $\pi_{\theta}(a_{t}\mid s_{t})$, where $\theta$ represents the parameters of the neural network. The policy maps observed features to a discrete set of agent behaviors, such as movement or object interaction.
This action is then executed in the environment, which returns a scalar reward $r_{t}$ and the next state $s_{t+1}$. In parallel, a prompt generation module extracts key contextual signals, such as the number of remaining orders and the current timestep, and formats them into a text prompt (e.g., "orders:2 t:25/400").
This textual prompt is evaluated by a frozen large language model (GPT-Neo 1.3B). The LLM produces logits for specific tokens such as "good" and "bad", which reflect the model’s confidence that the agents are cooperating effectively. If the logit for "good" exceeds that for "bad", a scalar bonus $\lambda$ is applied:
$\lambda=\begin{cases}\lambda_{\text{bonus}},&\text{if }\text{logit}(\texttt{"good"})>\text{logit}(\texttt{"bad"})\\ 0,&\text{otherwise}.\end{cases}$ (1)
If the LLM provides a positive evaluation, the agent’s reward becomes:
$r^{\prime}_{t}=r_{t}+\lambda.$ (2)
Otherwise, the original reward $r_{t}$ is retained. The shaping coefficient $\lambda$ is manually tuned to remain small and bounded, ensuring stability during PPO training.
The shaped reward $r^{\prime}_{t}$ is used to compute the advantage function $\hat{A}_{t}$, which guides PPO policy updates. We adopt the concept of generalized advantage estimation (GAE) [36], which links short-term reward shaping with long-horizon credit assignment. The temporal difference error is defined as:
$\delta_{t}=r^{\prime}_{t}+\gamma V(s_{t+1})-V(s_{t}).$ (3)
Here, $\gamma$ is the discount factor, and $V(s)$ is the learned value function for state $s$.
$\hat{A}_{t}=\delta_{t}+(\gamma\lambda)\delta_{t+1}+(\gamma\lambda)^{2}\delta_{t+2}+\cdots+(\gamma\lambda)^{T-t+1}\delta_{T-1}.$ (4)
Because the LLM remains frozen, its evaluations are deterministic and repeatable. These evaluations instantiate two fundamental game-theoretic principles. The first is belief revision, which parallels Bayesian updating: the LLM adjusts its cooperative judgment in response to evolving prompt content that reflects the current environment state, such as the number of pending orders, role distribution, or elapsed time. This mechanism allows the agent to revise expectations mid-task without retraining. The second principle is strategic consistency, which is implemented by reinforcing only those actions that align with collaborative success. Agents receive positive shaping signals only when their behavior contributes meaningfully to shared objectives, thereby discouraging selfish or unstable strategies and encouraging long-term alignment.
To operationalize these principles, we embed structured symbolic evaluation into the reinforcement learning loop. A prompt generation module continuously encodes task-relevant context, such as pending order count, current timestep, and relative agent positions, which is passed to a frozen LLM that produces a binary cooperative judgment. This symbolic signal is then mapped to a scalar bonus and integrated into each agent’s PPO training pipeline in real time. By providing feedback on the overall quality of collaboration, rather than low-level action success, the framework guides agents toward strategic behaviors that prioritize task flow, anticipation of partner actions, and recovery from errors. This allows for dynamic behavioral adjustments during execution, even in the presence of degraded perception or unpredictable task disruptions. As a result, agents adapt their policies mid-episode using high-level, interpretable feedback. This architecture bridges low-level reinforcement learning with symbolic reasoning, enabling robust coordination, adaptive role assignment, and resilience to uncertainty, including observation noise, partial observability, and real-time variability.
By linking symbolic judgment to policy adaptation, the framework naturally enforces game-theoretic alignment over time. Both PPO agents respond independently to the same symbolic signal and adapt their strategies to maximize cooperative feedback. As training progresses, the agents converge toward a stable configuration in which each agent’s role complements the other. This convergence satisfies the key condition of a Nash equilibrium between two agents: once both agents reach their equilibrium strategies, neither can improve their outcome by deviating unilaterally. Importantly, the framework does not solve for equilibrium analytically. Instead, equilibrium-like behavior emerges through iterative feedback and reward shaping, without any need for direct coordination or access to the co-agent’s internal state.
In summary, we have developed a new real-time shaping framework that integrates symbolic evaluation from frozen LLMs with policy optimization, enabling agents to adapt strategies dynamically in cooperative multi-agent settings. By fusing high-level symbolic reasoning with low-level reinforcement learning, our framework allows agents to revise behavior in real time without the need for retraining, message-passing, or centralized control. This approach offers several key advantages: (1) it provides interpretable, prompt-driven feedback that shapes coordination decisions at each step; (2) it enables agents to handle uncertainty, teammate variability, and perceptual noise through context-aware shaping; and (3) it facilitates convergence to equilibrium strategies via reward modulation rather than direct policy manipulation. The architecture grounds strategic behavior in game-theoretic principles such as belief revision and role alignment, supporting robust adaptation in dynamic environments. These properties make the system broadly applicable to settings where policies must remain flexible, resilient, and efficient under changing cooperative demands. We now evaluate its performance through a series of controlled simulations.
Figure 2: Custom Overcooked-AI “Burger Kitchen” scenario used in our experiments. Agent A (blue) retrieves raw patties from the fridge and cooks them on the Stove, while Agent B (green) assembles cooked patties with buns and condiments at the Burger Station and delivers the finished burger through the Delivery Window. Dashed arrows indicate task flow between stations. This scenario is a good illustration of our model because it imposes strict temporal dependencies and spatial constraints that require real-time role specialization, dynamic coordination, and mid-episode strategy adjustment. These are exactly the types of challenges our PPO+LLM shaping framework is designed to address.
3 Simulation Results and Analysis
We evaluate our framework in the Overcooked-AI simulator[35], a cooperative multi-agent benchmark in which two agents must complete cooking tasks under time pressure and partial observability. In our custom burger kitchen scenario (Figure 2), Agent A must wait while the patty cooks, then hand it off to Agent B for assembly and delivery. This layout enforces role specialization, timing dependencies, and dynamic coordination challenges that test the ability of our PPO+LLM system to adapt strategies based on real-time symbolic feedback.
As illustrated in Figure 2, the symbolic reward shaping framework is instantiated in this environment, where agents coordinate over spatially separated roles. By linking symbolic judgment to policy adaptation, the framework gradually enforces game-theoretic alignment. Both PPO agents respond independently to the same symbolic signal and adapt their behavior to maximize cooperative feedback. Over time, they converge toward a stable configuration in which each agent’s role complements the other.
This emergent structure satisfies the core property of a Nash equilibrium between two agents: once both adopt equilibrium strategies, neither can improve its outcome by deviating unilaterally. Importantly, the framework does not solve for equilibrium analytically. Instead, this behavior emerges through iterative learning guided by symbolic feedback, without requiring direct coordination or access to the co-agent’s internal state.
Figure 3: Mean return across clean and visibility noise environments. PPO+LLM performs comparably in clean settings and maintains advantage under perceptual uncertainty.
Figure 4: Mean return across noisy environments (Order Timing Noise and Combined Noise). PPO+LLM demonstrates improved robustness over PPO under multi-agent coordination disruptions.
We conduct a systematic evaluation of our PPO+LLM framework using the Overcooked-AI environment to analyze how LLM-guided reward shaping influences multi-agent collaboration, task efficiency, and robustness to uncertainty. Each model is tested over 200 independent episodes per environment, with controlled random seeds and standardized starting conditions to support consistent and fair evaluation.
We evaluate both standard PPO and PPO+LLM agents across four environments of increasing complexity: Clean (no noise), Visibility Noise, Order Timing Noise, and Combined Noise. The metrics include returns per episode, standard deviation, and average per-step latency. We report and analyze results across three core dimensions: performance, robustness, and inference efficiency.
Figure 5: Qualitative comparison in Overcooked-AI. Left: PPO+LLM agents coordinate by specializing in distinct roles and maintaining clear handoffs. Right: PPO-only agents often block each other, duplicate roles, or stall at delivery. The frozen LLM encourages efficient collaboration by rewarding high-level intent and coordination.
3.1 Returns per Episode
Figure 6: Burger Kitchen: Coordination Under Order Timing Noise. Left: PPO+LLM agents adapt to temporal perturbations and successfully coordinate task flow despite misaligned order signals. The tomato icon represents an off-timing ingredient request, which the symbolic feedback loop helps agents recover from. Right: PPO baseline agents fail to deliver the completed burger due to miscoordination, resulting in idle output and a missed delivery. This figure illustrates how symbolic feedback enables robust strategy adaptation under timing noise.
Figure 3 summarizes average returns per episode for PPO and PPO+LLM. In the Clean setting, where agents observe the environment without any disturbance, PPO fails to achieve any reward. This shows that the agent fails to develop cooperative strategies without step-by-step feedback guiding its actions during the episode. However, PPO+LLM consistently has a return of 9.8 with zero variance, demonstrating that binary LLM evaluations are sufficient to support reliable task completion.
Under the Visibility Noise condition, where the agent’s input state is interrupted with a random amount of perceptual noise, PPO again collapses to a return of 0. This outcome shows PPO’s sensitivity to low-level visual disturbances. In contrast, PPO+LLM stays strong and achieves a perfect return of 9.8 again. Since it learns from prompt-driven evaluations that do not rely on precise raw data, PPO+LLM remains stable even when perception is inconsistent.
The shaping mechanism becomes even more impactful under task-level disruption. In the Order Timing Noise setting, where task arrivals are randomly delayed or reordered, PPO performs poorly and has an average return of 40.4-40.4- 40.4 (standard deviation 3.7). This exposes a lack of stable learning which leads to low-quality actions. PPO+LLM significantly mitigates this degradation, recovering over 10 points (mean: 30.3-30.3- 30.3, std: 4.4). By consistently evaluating overall task progression (e.g., timestep and pending orders), the LLM provides structure that mitigates the unstable time dynamics that PPO alone fails to handle.
This pattern stays the same in the Combined Noise condition, where both visibility and timing are altered. PPO deteriorates to a mean of 39.9-39.9- 39.9 (std: 3.6), while PPO+LLM recovers nearly 25% of that loss with a mean return of 29.6-29.6- 29.6 (std: 4.3). This shows how symbolic feedback helps learning when rewards are weak or inconsistent, by giving the agent a clearer understanding of progress.
Figure 7: Burger Kitchen: Coordination Under Observation Noise. Left: PPO+LLM agents maintain effective collaboration despite partial visual occlusion. Even when perceptual details are masked or distorted, symbolic feedback helps agents preserve task flow and avoid conflicts. Right: PPO baseline agents fail to recognize delivery opportunities due to limited visibility, leading to idle burgers and missed outputs. This figure demonstrates how language-based shaping enables perceptual robustness and consistent goal pursuit under noisy observations.
To illustrate how the system works in practice, we present a qualitative example from the Overcooked-AI environment (Figure 5). In this episode, one agent retrieves ingredients while the other prepares plates and manages deliveries. The frozen LLM rewards states where agents remain active, coordinate handoffs efficiently, and avoid interfering with one another. This leads to smooth, high-throughput cooperation. In contrast, the PPO-only baseline often results in role duplication, hesitation, or spatial conflicts between agents. This example shows that symbolic feedback improves not just numerical performance but also the structure and clarity of collaborative behavior.
Figure 8: Burger Kitchen: Coordination Under Combined Noise. Left: PPO + LLM agents handle both mis-timed ingredient requests (clock icon) and sensor corruption (pixel-noise tile), reassigning roles on the fly and completing delivery. Right: PPO baseline agents break the task pipeline; perception noise triggers hesitation, timing jitter causes the burger to arrive too late, and the meal sits idle ($\times$) at the window. The comparison shows how prompt-based symbolic feedback enables recovery from multiple simultaneous disturbances.
To evaluate coordination performance under temporal uncertainty, we test both PPO and PPO+LLM agents in the Order Timing Noise environment, where requested ingredients arrive at irregular intervals. This setting introduces misaligned task dependencies, increasing the risk of agents acting prematurely or stalling due to incomplete orders. As shown in Figure 6, PPO+LLM agents recover from off-cycle prompts (e.g., a tomato requested ahead of time) by dynamically reassigning roles and maintaining structured task flow. In contrast, PPO baseline agents lack a mechanism to reconcile timing inconsistencies, leading to delivery failures and idle outputs. This scenario highlights the effectiveness of prompt-based symbolic guidance in helping agents reason through dynamic task adjustments, enabling more stable coordination even when order streams are perturbed.
To evaluate robustness under perceptual uncertainty, we test agents in the Observation Noise environment, where input states are partially masked or corrupted prior to reaching the policy network. This simulates scenarios in which agents must act with incomplete or degraded visibility. As shown in Figure 7, PPO+LLM agents successfully maintain cooperative behavior, relying on symbolic evaluations that remain stable across noisy visual inputs. The frozen LLM reinforces high-level intent rather than low-level pixel accuracy, enabling agents to avoid collisions, infer teammate roles, and complete deliveries even when specific object placements are obscured. In contrast, PPO baseline agents often misinterpret the scene, leading to duplicated efforts or missed handoffs. This setting highlights the value of prompt-based feedback in providing consistent behavioral signals, especially when raw observations are unreliable.
To evaluate agent performance under simultaneous disturbances, we consider the Combined Noise environment, which introduces both observation corruption and task timing irregularities. This dual-source disruption challenges agents to interpret partially noisy visual inputs while also adjusting to asynchronous order arrivals. As shown in Figure 8, PPO baseline agents frequently fail in this setting. Visual distortion causes hesitation or miscoordination, and mistimed ingredient requests result in broken task flow, leading to idle or undelivered meals. In contrast, PPO+LLM agents use symbolic feedback to remain adaptive. The LLM reinforces cooperative behavior even under perceptual ambiguity, enabling agents to reassess roles and recover from misalignments. This robustness allows PPO+LLM to maintain productive collaboration despite multiple overlapping sources of uncertainty, demonstrating the effectiveness of language-based shaping in complex, real-time environments.
3.2 Robustness to Noise
Figure 4 shows the average return per episode under task-level noise. We evaluate each method’s sensitivity to structured disturbances to determine its robustness. To simulate insufficient state visibility, visibility noise modifies the grid-level tensor prior to it reaching the policy network. PPO collapses under noisy conditions due to its dependence on unal tered input encodings for fine-grained value learning, resulting in no reward accumulation. In contrast, PPO+LLM maintains positive returns across both conditions. This demonstrates that symbolic evaluations from the LLM offer resilience by reinforcing cooperative intent, even when perception is corrupted or incomplete.
Since PPO+LLM’s prompt-based evaluation relies on higher-level elements like order count and timestep, which are not directly harmed by visual corruption, it remains unaffected and keeps an average return of 9.8.
Order Timing Noise introduces uncertainty into the environment’s causal structure by changing when new objectives are presented. This disrupts the routine coordination patterns PPO often overfits to. Without the ability to anticipate task sequences, PPO’s policy collapses. PPO+LLM, however, keeps its strategic framework intact by relying on ongoing LLM-guided assessments to judge whether progress is being made. Its behavior remains adaptive, shaped by a reward signal that is context-aware and flexible.
Combined Noise, which is arguably the most difficult condition, introduces both perceptual distortion and goal-timing irregularity, disrupting low-level coordination patterns. PPO’s performance deteriorates substantially in this setting. In contrast, PPO+LLM continues to produce reliably higher returns by leveraging symbolic feedback from LLM prompts. As shown in Figure 4, this feedback serves as an external corrective mechanism, helping agents focus on high-level cooperation rather than brittle stimulus-response loops. This enables more resilient coordination even when the environment becomes unpredictable or misaligned with learned routines.
Figure 9: Latency per decision step across environments. PPO+LLM adds modest computational cost but stays within real-time limits across all conditions.
3.3 Latency Tradeoff
Figure 9 summarizes the latency per decision step across four environments. While PPO+LLM delivers improvements in robustness and strategic coordination, it introduces a lightweight runtime cost due to real-time prompt generation and LLM-guided evaluation. In the Clean environment, PPO’s per-step latency averages 0.77 milliseconds, whereas PPO+LLM increases this to 0.99 milliseconds, introducing an approximately 28% overhead caused by token inference and string formatting. Similar latency increases are observed across all settings, but PPO+LLM consistently remains below 1.05 milliseconds per step, ensuring that performance remains within real-time constraints.
Importantly, this overhead is both small and stable. Despite task randomness and structured disturbances, the shaping module’s runtime does not fluctuate due to its reliance on a frozen LLM and a deterministic prompt structure. This predictability is critical for downstream integration in safety-sensitive or latency-bound applications. In addition, because the LLM is queried with compact, template-based prompts, the shaping mechanism remains computationally efficient and scales well across multi-agent deployments. These properties make PPO+LLM viable not just for simulated benchmarks, but also for interactive robotics, collaborative agents, and fast-paced games where decision-time guarantees are essential.
Figure 10: Task-Completion Rate (9.4$\geq$ 9.4≥ 9.4) across environments. PPO+LLM reliably completes full tasks in clean and perceptually noisy settings, while PPO consistently fails to reach the success threshold.
3.4 Task Completion Rate
To evaluate whether agents successfully completed the full collaborative task, we introduced a completion-rate metric that counts the percentage of episodes where the total return exceeded a threshold of 9.4. This cutoff was chosen based on empirical observations, where PPO+LLM agents consistently reached scores near 9.8 in successful episodes. Setting the threshold slightly below 9.8 allows for small variations while still representing full task completion.
As shown in Figure 10, PPO+LLM agents achieved a 100% completion rate in both the clean and observation noise settings, while PPO agents consistently failed to meet the threshold in all environments. In more challenging conditions, such as order timing noise and combined noise, neither policy consistently completed the task, resulting in 0% success. However, PPO+LLM still demonstrated substantially higher average returns in these noisy environments, indicating partial progress toward task completion despite increased difficulty.
These results suggest that LLM-guided agents are more capable of executing full task cycles in favorable or moderately noisy environments and maintain more robust performance under degraded conditions, even when full completion is not achieved. To qualitatively demonstrate this coordination, next we visualize the behavioral progression toward equilibrium in a concrete kitchen setting.
Figure 11: Convergence to Nash equilibrium in a 2-agent Burger Kitchen task. PPO-LLM agents transition from initial miscoordination (Step 1) to stable cooperative behavior (Step 4), resulting in a mutually optimal joint policy where neither agent benefits from deviating.
Figure 11 illustrates how PPO+LLM agents converge toward a Nash equilibrium in the Burger Kitchen environment. In Step 1, the agents operate with no coordination, as both attempt to cook or deliver, causing collisions and idle time. In Step 2, one agent begins to approach the burger station while the other still overlaps tasks, leading to partial role confusion. By Step 3, the agents implicitly specialize: one retrieves and cooks patties, while the other handles burger assembly and delivery. Finally, in Step 4, the agents settle into complementary roles that form a stable cooperative policy. Agent A consistently handles cooking, while Agent B manages final assembly and delivery. This strategy consistently achieves high episodic returns (average 9.8 out of 10) and full task completion.
Importantly, this configuration forms a Nash equilibrium because unilateral deviation results in lower total rewards. For example, when Agent A deviates by moving toward the delivery station instead of cooking, both agents experience bottlenecks and blocking, reducing the return to an average of 6.2. Similarly, when Agent B attempts to cook instead of delivering, assembled burgers are delayed or missed entirely, leading to an average reward of 5.9. Since any deviation by one agent reduces the collective outcome and does not benefit the deviating agent individually, both agents have no incentive to alter their strategy once the cooperative equilibrium is reached. This visual progression demonstrates that language-guided agents can internalize equilibrium strategies in dynamic, multi-agent environments and adapt toward optimal joint behavior.
4 Conclusion and Future Work
In this paper, we have proposed a framework that integrates frozen large language models with reinforcement learning agents to support real-time strategy shaping in cooperative multi-agent environments. We have introduced symbolic feedback in the form of prompt-based binary evaluations, which refined low-level decision-making without requiring LLM fine-tuning or online optimization. We have evaluated the system across four increasingly noisy Overcooked-AI environments and found that PPO+LLM agents consistently outperformed PPO baselines. In high-noise conditions, we have achieved up to a 26% improvement in average return, while maintaining inference latency below 1.05 milliseconds per decision step. While our framework supports real-time shaping across multiple decision steps, it does not incorporate dialogue-based interactions, such as conversational feedback from the LLM or language-mediated coordination between agents, which may offer richer strategic reasoning and alignment. These forms of interactive feedback and inter-agent communication represent promising directions for future extension, potentially enabling more adaptive and socially aware multi-agent behavior.
Beyond these extensions, several additional research directions remain open. Our current system treats the LLM as a one-shot evaluator. However, enabling iterative and bidirectional interaction between agents and LLMs could enhance strategic grounding and help disambiguate complex decisions. Future frameworks may also incorporate agent-to-agent communication via natural language, supporting joint planning, intent signaling, and belief sharing. While our agents receive symbolic shaping within an episode, they currently do not persist or generalize feedback across episodes. Introducing memory modules or belief tracking could facilitate long-term strategy formation. Finally, although Overcooked-AI serves as a structured benchmark, future research should investigate more complex domains such as competitive negotiation, open-ended roleplay, or real-time robotics. These settings require symbolic reasoning, adaptive coordination, and language-guided shaping, and will further advance the development of autonomous, socially aware, and interpretable multi-agent systems.